OpenStudy (anonymous):

simplify: (100/49)^(-1/2)

OpenStudy (anonymous):

Google it...

OpenStudy (anonymous):

For removing negative sign, just switch denominator and numerator.. Like: \[\large (\frac{x}{y})^{-a} \implies \color{blue}{(\frac{y}{x})^{a}}\]

OpenStudy (anonymous):

So what will it become by removing negative sign: \[\huge (\frac{100}{49})^{\frac{-1}{2}} = ??\]

OpenStudy (anonymous):

(100/49)^(-1/2) = 1/(100/49)^(1/2)

OpenStudy (anonymous):

\[1/\sqrt{100/49}\] = 1/(10/7)

OpenStudy (anonymous):

Then reverse it too..

OpenStudy (anonymous):

1/1 x 7/10 = 7/10

OpenStudy (anonymous):

Yeah...that's what google said, too... :-)

OpenStudy (anonymous):

\[\large (\frac{100}{49})^{\frac{-1}{2}} \implies (\frac{49}{100})^{\frac{1}{2}} \implies \sqrt{\frac{49}{100}} \implies ??\]